# From quantum enhanced to quantum inspired Monte Carlo

### v1

@@ Appendix B Bayesian optimization of time-dependent schedules
In Bayesian optimization, the goal is to maximize a black-box function $f$ defined over a parameter space $D_{\theta}\subset\mathbb{R}^{m}$. BO constructs a surrogate model to approximate the objective function, often using Gaussian Processes (GPs). The GP model, $\tilde{f}(\theta)$, is characterized by a mean function $\mu(\theta)$ and a kernel function $k(\theta,\theta^{\prime})$, which captures the correlation between points. Here, we employ the Matérn 5/2 kernel [77]. By conditioning the GP on observed data, the model updates its predictions, incorporating new information. The updated GP is then used to determine the next point to probe by maximizing the acquisition function $\phi(\theta;\kappa)$. Here, we use the Upper Confidence Bound (UCB), defined as:
@@
where $\sigma(\theta)$ represents the predicted uncertainty. The acquisition function balances exploration (sampling regions with high uncertainty) and exploitation (sampling regions with high predicted values). As the optimization progresses, the parameter $\kappa$ is decreased to gradually shift the focus from exploration to exploitation. This iterative process of updating the surrogate model and selecting new points based on the acquisition function continues until convergence. BO has recently been employed to design quantum annealing schedules [50] and to optimize the parameters of the quantum approximate optimization algorithm (QAOA) [78].
In our case, the objective is to optimize the shape of $\gamma(t)$ to produce a proposal distribution with the maximum absolute spectral gap. We fix the total evolution time at $\tau=10$ and define the dimensionless fraction $s(t)=\frac{t}{\tau}$ of the total evolution time. Initially, $\gamma(s(t))$ is parameterized by selecting 5 equidistant points. Intermediate values are then obtained using cubic interpolation, resulting in a piecewise cubic schedule. Additionally, we enforce $\gamma(0)=\gamma(\tau)=0$ to ensure that the schedule begins with a classical Hamiltonian, thereby allowing the perturbation to be switched on gradually.
The cost function in our optimization problem takes a 5-dimensional vector $\theta$, evaluates the proposal distribution by integrating the time-dependent Schrödinger equation, and calculates the absolute spectral gap of the transition matrix, which serves as our figure of merit.
We perform the aforementioned optimization for system sizes ranging from $n=4$ to $n=9$. The resulting optimized schedules are depicted in Fig. 12. Remarkably, we observe that the optimizer converges to similar schedule shapes for system sizes $n\in[4,5]$, $n\in[6,7]$, and $n\in[8,9]$ respectively. For $n\in[6,7]$, the schedule ramps up the transverse field sharply to $\gamma\approx 1$, where the Hamiltonian is dominated by the mixing term. In contrast, for $n\in[8,9]$, the optimized schedule resembles the time-independent case. Here, $\gamma(t)$ increases to approximately its critical value, maintains an almost constant field strength, and then symmetrically decreases towards the end.
Figure 12: Optimized schedules $\gamma(s)$ for different system sizes ($n=4$ to $n=9$) obtained using BO. The x-axis represents the dimensionless fraction $s=\frac{t}{\tau}$. The schedules are constrained to be symmetric around the middle point. The critical value $\gamma_{c}$ is indicated for reference.
@@ Appendix C Violation of Detailed Balance
The trotter error does not impact the symmetry condition, because the $U$ and its transpose $U^{T}$ only differ by a phase. We can verify this by expressing the 2nd order trotterized unitary in terms of its transpose and by noting that the $e^{-iH_{Z}}$ is a diagonal matrix in the computational basis:
$U(t)_{\text{2nd}}=e^{-2iH_{Z}dt}(U(t)_{\text{2nd}})^{T}e^{+2iH_{Z}dt}=(U(t)_{\text{2nd}})^{T}$ (25)
$U(t)_{\text{2nd}}$ is defined in equation 26. In the verification, we used the fact that $H_{Z}=H_{Z}^{T}$ and $H_{X}=H_{X}^{T}$ are symmetric and that for matrices $A,B,C$ the following identities hold: $(ABC)^{T}=C^{T}B^{T}A^{T}$, $(A^{n})^{T}=(A^{T})^{n}$, $(A+B)^{T}=A^{T}+B^{T}$ and therefore $(e^{A})^{T}=e^{(A^{T})}$.
@@ Appendix D Impact of approximation errors on the scaling
Figure  9 in subsection  V.2.2 depicts the scaling of the spectral gap in system size for different proposal strategies at fixed temperature $T=1$. Figure  13 depicts the scaling of the spectral gap in temperature $T$ for different proposal strategies at fixed system size $n=8$. We observe that the performance of the MPS proposals varies highly with the bond dimension $\chi$.
Figure 13: Spectral gap $\delta$ for different temperatures $T$ at fixed system size $n=8$ for different proposal strategies. $\gamma=0.45$, $t=12$. Trotter $dt=0.8$ refers to the trotterized Quantum proposal. MPS $\chi$ refers to using TEBD with MPS to perform the time evolution with bond dimension $\chi$. Each data point is averaged over 100 instances. $\chi=16$ would result in the exact representation of the quantum trotter proposal.
@@ Appendix E Details of the Tensor-network calculations
This Section details the implementation of the quantum-inspired proposal step in Sect. V.2. The numerical simulations are implemented in the Julia Programming language using the $iTensor.jl$ library [68]. The proposal requires us to implement the $|\psi(t)\rangle=U|\psi(0)\rangle$. Given the Hamiltonian $H=(1-\gamma)\alpha H_{c}+\gamma H_{mix}=H_{Z}+H_{X}$, we split $U(t)=e^{-iHt}$ into the following product
$U(t)_{\text{2nd Order }}=e^{+iH_{Z}\frac{dt}{2}}\left[U(t)_{1\text{ stOrder }}\right]e^{-iH_{Z}\frac{dt}{2}}$ (26)
to get the 2nd order trotterized unitary where
$U(t)_{1\text{ st Order }}=e^{-iH_{z}^{(1)}dt}e^{-iH_{x}^{(1)}dt}\ldots e^{-iH_{z}^{(m)}dt}e^{-iH_{X}^{(m)}dt}$
with $t=mdt$. The first and second-order trotterized unitary only differ by a phase, which is irrelevant for the associated probability distribution.
We further divide trotterized unitary into one-site and two-site gates. Assuming a complete (or fully connected) graph of $n$ spins, there are $n$ vertices and $n(n-1)/2$ edges. Therefore, we have to construct $n$ $Z$-gates, $n$ $X$-gates and $n(n-1)/2$ ZZ-gates. The $X$ and $Z$ gates are single-site gates of the form
$G\left(Z_{\sigma_{j}}\right)=e^{-idt\gamma Z_{\sigma_{j}}};\quad G\left(X_{\sigma_{j}}\right)=e^{-idt(1-\gamma)X_{\sigma_{j}}}$
and are represented as rank-2 tensors ( 2 indices) where each index is of dimension $d=2$ for the spin-$1/2$ particles. The $ZZ$-gates are of the form
@@
and are represented as rank-4 tensors where each index is of dimension $d=2$ as well. The one-site and two-site gates are the building blocks required for the Time Evolving Block Decimation (TEBD) of the Matrix Product States (MPS). The procedure can be separated into the following steps:
  1. 1.
@@
    System Setup: The underlying (complete) graph of spins and their tensor indices are created. The graph is mapped to an MPS. The MPS is initialized.
  2. 2.
@@
    Gate Construction: The Unitary evolution is trotterized and split into one-site gates (related to the X or Z terms in Hamiltonian) and two-site gates (related to the ZZ terms in Hamiltonian). The gates represent an evolution step and are computed as matrix exponentials of the Hamiltonian.
  3. 3.
@@
    Gate Application: The time evolution of the state is done by applying the gates to the MPS.
  4. 4.
    Normalization: After each time step the MPS is normalized.
  5. 5.
@@
    Sampling: The state is sampled after concluding the time evolution, which results in the new proposal step.

### v2

@@ Appendix B Bayesian optimization of time-dependent schedules
In Bayesian optimization, the goal is to maximize a black-box function $f$ defined over a parameter space $D_{\theta}\subset\mathbb{R}^{m}$. Bayesian optimization constructs a surrogate model to approximate the objective function, often using Gaussian processes (GPs). The GP model $\tilde{f}(\theta)$ is characterized by a mean function $\mu(\theta)$ and a kernel function $k(\theta,\theta^{\prime})$, which captures the correlation between points. Here we employ the Matérn 5/2 kernel [78]. By conditioning the GP on observed data, the model updates its predictions, incorporating new information. The updated GP is then used to determine the next point to probe by maximizing the acquisition function $\phi(\theta;\kappa)$. Here we use the upper confidence bound, defined as
@@
where $\sigma(\theta)$ represents the predicted uncertainty. The acquisition function balances exploration (sampling regions with high uncertainty) and exploitation (sampling regions with high predicted values). As the optimization progresses, the parameter $\kappa$ is decreased to gradually shift the focus from exploration to exploitation. This iterative process of updating the surrogate model and selecting new points based on the acquisition function continues until convergence. Bayesian optimization has recently been employed to design quantum annealing schedules [49] and to optimize the parameters of the QAOA [79].
In our case, the objective is to optimize the shape of $\gamma(t)$ to produce a proposal distribution with the maximum absolute spectral gap. We fix the total evolution time at $\tau=10$ and define the dimensionless fraction $s(t)=\frac{t}{\tau}$ of the total evolution time. Initially, $\gamma(s(t))$ is parametrized by selecting five equidistant points. Intermediate values are then obtained using cubic interpolation, resulting in a piecewise cubic schedule. Additionally, we enforce $\gamma(0)=\gamma(\tau)=0$ to ensure that the schedule begins with a classical Hamiltonian, thereby allowing the perturbation to be switched on gradually.
The cost function in our optimization problem takes a five-dimensional vector $\theta$, evaluates the proposal distribution by integrating the time-dependent Schrödinger equation, and calculates the absolute spectral gap of the transition matrix, which serves as our figure of merit.
We perform the aforementioned optimization for system sizes ranging from $n=4$ to $9$. The resulting optimized schedules are depicted in Fig. 12. Remarkably, we observe that the optimizer converges to similar schedule shapes for system sizes $n\in[4,5]$, $n\in[6,7]$, and $n\in[8,9]$, respectively. For $n\in[6,7]$, the schedule ramps up the transverse field sharply to $\gamma\approx 1$, where the Hamiltonian is dominated by the mixing term. In contrast, for $n\in[8,9]$, the optimized schedule resembles the time-independent case. Here $\gamma(t)$ increases to approximately its critical value, maintains an almost constant field strength, and then symmetrically decreases towards the end.
Figure 12: Optimized schedules $\gamma(s)$ for different system sizes obtained using BO: (a) $n=4$, (b) $n=5$, (c) $n=6$, (d) $n=7$, (e) $n=8$, and (f) $n=9$.The x axis represents the dimensionless fraction $s=\frac{t}{\tau}$. The schedules are constrained to be symmetric around the middle point. The critical value $\gamma_{c}$ is indicated for reference.
@@ Appendix C Violation of Detailed Balance
The Trotter error does not impact the symmetry condition, because the $U$ and its transpose $U^{T}$ only differ by a phase. We can verify this by expressing the second-order Trotterized unitary in terms of its transpose and by noting that the $e^{-iH_{Z}}$ is a diagonal matrix in the computational basis:
$U(t)_{\text{2nd}}=e^{-2iH_{Z}dt}[U(t)_{\text{2nd}}]^{T}e^{+2iH_{Z}dt}=[U(t)_{\text{2nd}}]^{T}$ (25)
The $U(t)_{\text{2nd}}$ is defined in Eq. 26. In the verification, we used the fact that $H_{Z}=H_{Z}^{T}$ and $H_{X}=H_{X}^{T}$ are symmetric and that for matrices $A,B$ and $C$ the identities $(ABC)^{T}=C^{T}B^{T}A^{T}$, $(A^{n})^{T}=(A^{T})^{n}$, and $(A+B)^{T}=A^{T}+B^{T}$ hold and therefore $(e^{A})^{T}=e^{(A^{T})}$.
@@ Appendix D Impact of approximation errors on the scaling
Figure  9 in Sec. V.2.2 depicts the scaling of the spectral gap in system size for different proposal strategies at fixed temperature $T=1$. Figure  13 depicts the scaling of the spectral gap in temperature $T$ for different proposal strategies at fixed system size $n=8$. We observe that the performance of the MPS proposals varies greatly with the bond dimension $\chi$.
Figure 13: Spectral gap $\delta$ for different temperatures $T$ at fixed system size $n=8$ for different proposal strategies, with $\gamma=0.45$ and $t=12$. Trotter $dt=0.8$ refers to the Trotterized quantum proposal. MPS $\chi$ refers to using TEBD with the MPS to perform the time evolution with bond dimension $\chi$. Each data point is averaged over 100 instances. Bond dimension $\chi=16$ would result in the exact representation of the quantum Trotter proposal.
@@ Appendix E Details of the Tensor-network calculations
This appendix details the implementation of the quantum-inspired proposal step in Sec. V.2. The numerical simulations are implemented in the Julia programming language using the itensor.jl library [68]. The proposal requires us to implement the $|\psi(t)\rangle=U|\psi(0)\rangle$. Given the Hamiltonian $H=(1-\gamma)\alpha H_{c}+\gamma H_{mix}=H_{Z}+H_{X}$, we split $U(t)=e^{-iHt}$ into the product
$U(t)_{\text{2nd order }}=e^{+iH_{Z}(dt/2)}\left[U(t)_{1\text{st order}}\right]e^{-iH_{Z}(dt/2)}$ (26)
to get the second-order Trotterized unitary where
$U(t)_{\text{1st order}}=e^{-iH_{z}^{(1)}dt}e^{-iH_{x}^{(1)}dt}\cdots e^{-iH_{z}^{(m)}dt}e^{-iH_{X}^{(m)}dt},$
with $t=mdt$. The first and second-order Trotterized unitaries only differ by a phase, which is irrelevant for the associated probability distribution.
We further divide the Trotterized unitaries into one-site and two-site gates. Assuming a complete (or fully connected) graph of $n$ spins, there are $n$ vertices and $n(n-1)/2$ edges. Therefore, we have to construct $n$ $Z$ gates, $n$ $X$ gates, and $n(n-1)/2$ $ZZ$ gates. The $X$ and $Z$ gates are single-site gates of the form
$G\left(Z_{\sigma_{j}}\right)=e^{-idt\gamma Z_{\sigma_{j}}},\quad G\left(X_{\sigma_{j}}\right)=e^{-idt(1-\gamma)X_{\sigma_{j}}}$
and are represented as rank-2 tensors (two indices) where each index is of dimension $d=2$ for the spin-$1/2$ particles. The $ZZ$ gates are of the form
@@
and are represented as rank-4 tensors where each index is of dimension $d=2$ as well. The one-site and two-site gates are the building blocks required for the TEBD of the MPSs. The procedure can be separated into the following steps:
  1. (i)
@@
    System setup. The underlying (complete) graph of spins and their tensor indices is created. The graph is mapped to an MPS. The MPS is initialized.
  2. (ii)
@@
    Gate construction. The unitary evolution is Trotterized and split into one-site gates (related to the $X$ or $Z$ terms in the Hamiltonian) and two-site gates (related to the $ZZ$ terms in the Hamiltonian). The gates represent an evolution step and are computed as matrix exponentials of the Hamiltonian.
  3. (iii)
@@
    Gate application. The time evolution of the state is done by applying the gates to the MPS.
  4. (iv)
    Normalization. After each time step the MPS is normalized.
  5. (v)
@@
    Sampling. The state is sampled after concluding the time evolution, which results in the new proposal step.
Appendix F Memory and Computational Cost Estimates of TEBD for MPS
This appendix details the assumptions and derivations in the memory and computational cost estimates for the local, uniform, and MPS proposal strategies. The estimates serve as a basis for estimates of the time needed to perform an individual Markov chain step [$t_{c}(n)$ for local and uniform, and $t_{qi}$ for the QIMC] in Sec. V.2.3.
The local proposal strategy consists of randomly drawing one of $n$ spins and flipping it. This corresponds to storing $n$ complex coefficients and a computational complexity of $O(1)$. The uniform proposal also has to store the $n$ coefficients but draws $n$ spins randomly, resulting in a computational complexity of $O(n)$. Compared to the local and uniform proposal strategies, the MPS proposal strategy requires vastly more time and memory resources. In the following, we use $d$ for the dimension of the local spins, assume a fully connected spin graph with $n$ vertices and $n(n-1)/2$ edges, and look at the TEBD of the MPS as described in Appendix E.
The dominating memory costs are the storage of the matrix product state with $O(nd\chi^{2})$ and the storage of the application gates. We need to store $2n$ single-site gates (for the $X$ and $Z$ terms in the Hamiltonian) and $n(n-1)/2\propto n^{2}$ two-site gates (for the $ZZ$ terms in the Hamiltonian). The two-site gates are the dominating term and are stored as rank-4 tensors where each index is of dimension $d$ resulting in an $O(n^{2}d^{4})$ contribution. Therefore, the memory cost can be approximated by $O\left(nd\chi^{2}+n^{2}d^{4}\right)$. For instance, using 16 GB of memory, we can store an MPS and the associated gates for $n=16$ with the exact bond dimension of $\chi=256$ or $n=100$ with $\chi=95$. In both cases, we use complex coefficients and the single-precision floating-point format (float32).
In terms of computational time, we can neglect the system setup, the gate construction, the one-site gate applications, the normalization, and the sampling, because the dominating factor is the two-site gate application.
The gate application represents the time evolution of the state. For each Trotter step, we have to apply all one- and two-site gates once. In total, there are $m$ Trotter steps resulting in $t=mdt$. The application of a one-site gate can be done without incurring an error and without increasing the memory required. The number of contractions needed to apply the one-site gate to the MPS is $O\left(d^{2}\chi^{2}\right)$.
The application of a two-site gate is more complicated. We have to distinguish between sites that are adjacent in the MPS and sites that are non adjacent. For adjacent sites, the gate application consists of contracting all indices in an efficient order, reshaping the indices, applying a singular value decomposition (SVD), and performing a truncation to the desired bond dimension or accuracy. The contraction of all indices requires $O\left(d^{4}\chi^{2}+\chi^{3}d^{3}\right)=O\left(\chi^{3}d^{3}\right)$ operations. The SVD is performed in $O\left(\chi^{3}d^{3}\right)$ operations. If we do not truncate the bond dimension between the two sites, the gate application grows the bond dimension from $\chi$ to $\chi d$.
If the sites are not adjacent, there are different ways of applying the two-site gate. In the following, we assume that we have two sites $i$ and $j$ that are separated by $L=j-i+1$. The brute-force approach would require us to contract all intermediate bond indices between sites $i$ and $j$, reshuffle the indices, apply the two-site gate and decompose the resulting tensor back into MPS form. This is really inefficient.
To get a better estimate for the number of contractions necessary to perform a non adjacent two-site gate we follow the arguments provided in [73, 67] for long-range interactions in one-dimensional systems. The underlying idea is to use swap gates that shuffle the indices in the MPS. This encompasses swapping indices to create adjacent sites, applying the gate, and then swapping back the indices to their original positions. If the bond dimensions are not truncated, the bond dimensions between the two sites increase by $d$ or $d^{2}$.
Reshuffling the indices of a rank-$a$ tensor, where each index is of dimension $b$, requires $O(b^{a})$ operations. Reshuffling the physical indices of two adjacent MPS sites requires us to first contract the two adjacent indices to create a combined tensor $O(d^{2}\chi^{3})$, then reshuffle the indices of the new tensor $O(d^{2}\chi^{2})$, and finally apply an SVD to get back to the MPS form $O(d^{3}\chi^{2})$. Overall, one swap gate requires $O(d^{2}\chi^{3}+d^{2}\chi^{2}+d^{3}\chi^{3})=O(d^{3}\chi^{3})$ operations.
Different two-site gates require different numbers of swap gates depending on their adjacency. Out of the $n(n-1)/2$ two-site gates, $n-1$ require zero swap gates, $n-2$ require two swap gates, $n-3$ require four swap gates ending with one gate requiring $n-2$ swap gates. Hence, we have to perform a total of
$2\sum_{a=1}^{n-1}(n-a)(a-1)=\frac{n^{3}}{3}+n^{2}+\frac{2}{3}n$ (27)
swap operations for one Trotter step. Each swap gate is $O(d^{3}\chi^{3})$. Therefore, applying all two-site gates for one Trotter step requires operations of
$O\left(\left(\frac{n^{3}}{3}+n^{2}+\frac{2}{3}n\right)d^{3}\chi^{3}\right)=O\left(n^{3}d^{3}\chi^{3}\right),$ (28)
which is the largest contributing factor in the application of the two-site gate application. In addition, if we do not truncate the bond dimension after each gate application, the maximum bond dimension $\chi=2^{n/2}$ is reached after applying the first $n$ two-site gates, i.e., the two-site gates connecting the leftmost index with all the other indices. Accounting for the $m$ Trotter steps, this results in contractions of $O(mn^{3}d^{3}\chi^{3})$, as summarized in Table 1.